Every rational number can be expressed as a quotient p/q so that p and q has no common factors.


1
Expert's answer
2021-01-14T16:28:29-0500

Let the rational number be "a=\frac{p}{q}" , such that "p,q\in\Z" and gcd("p,q")=1


Let us assume that gcd("p,q")"\neq1" or gcd("p,q")=c(let), such that c"\in\Z"


Then we write a as:-



"a=\frac{p}{q}=\frac{\frac{p}{c}}{\frac{q}{c}}=\frac{p'}{q'}"

where, "cp'=p" and "cq'=q" , where "p',q'\in\Z" as "p,q,c\in\Z"


Then,



"\Rightarrow a=\frac{p'}{q'}"

where, "p',q'\in\Z" and gcd("p',q'")=1


Therefore, our assumption that gcd("p,q")=c still holds the definition of rational number.